# On the phase of the de Sitter density of states

### v1

@@ 1 Introduction
This puzzle was addressed by Maldacena in [7], motivated in part by the usefulness of including an observer in other studies of de Sitter space [8, 9]. [^2] It would be interesting to connect this proposal to the bulk vs. edge mode discussion in [10, 11]. The idea of [7] is that if an observer is included, all but two of the $D+1$ conformal Killing vectors become physical modes, because they move the observer. Or, in a different gauge, the observer adds D1D-1 new unstable modes that require Wick rotation. Either way, the phase now becomes $\mathrm{i}^{D+2-(D-1)}=\mathrm{i}^{3}=-\mathrm{i}$. [^3] Further work studied the phase for product manifolds [12, 13] and the Coleman-de Luccia instanton [14].
@@ 3.1 Setting up the computation
In the one-loop determinant, each negative mode gives a factor of $(-\mathrm{i})$, see appendix A. Our starting point in analyzing the phase is to define the measure so that it is real and positive if we Wick-rotate the entire conformal factor, or more generally one whole field [4]. In other words, we multiply the naive one-loop determinant by $(+\mathrm{i})^{n}$, where $n$ is the number of spacetime points. To evaluate the net phase, we would like to compare the number of negative eigenvalues of $\mathcal{D}$ to the number of spacetime points. In section 3.2, we use a pragmatic approach [^7] One could instead simply rotate one or the other of $\phi,\omega$ (or a linear combination) of them, but this leads to a complex fluctuation operator $\mathcal{D}$ that we found difficult to analyze. that allows this comparison to be made. We discretize the manifold with a mesh, on which we study the spectrum of the quadratic fluctuations. For a finite mesh, the number of negative eigenvalues of $\mathcal{D}$ and the number of spacetime points are both finite and can be compared. We check this numerical answer with an analytical computation in the ultracold limit and an analytical understanding of how the phase changes at the special point $Q=1/4$.
@@ 3.3 A model for the lukewarm black holes
It is interesting to compare to the computation of the Euclidean Schwarzschild black hole in 4D flat space. This has a negative mode in the canonical ensemble [34, 35, 17] that can be explained by a similar integral, over the area-radius of the horizon: [^11] See appendix F.1 of [32] and more recently [36] for a similar discussion of small black holes in AdS.
@@ 3.4 Ultracold Nariai
Here, $\omega_{\ell,m}$ are the spherical harmonic modes normalized using the physical metric $g$, not the metric $\widetilde{g}$. They are all unstable except the three $\ell=1$ zero modes, corresponding to the conformal Killing modes on the sphere (see (3.4)). The phase of the partition function comes from rotating all the unstable modes, each one gives a factor of $\mathrm{i}$. Due to the three missing zero modes, that leads to $\mathrm{i}^{3}$. The three zero modes also lead to a $q^{3}$ one-loop determinant. Together, these lead to the one-loop factor:
@@ 4 The phase of the state-counting partition function
In the previous sections, we studied the phase of the Euclidean gravity path integral. In this section, we compute the phase of the state-counting partition function. The difference is that the Euclidean gravity path integral is over real Euclidean geometries (up to contour rotations for convergence), whereas for the state-counting partition function we modify the integral in order to impose the Hamiltonian constraint $H=0$ [15, 16, 37, 38, 7].
@@ 4.3 On the overall sign
We suggest a different prescription, based on the idea that the $\beta$ integral is supposed to impose the Hamiltonian constraint. To implement this, it is necessary to expose at least one further integral, over the energy or area variable conjugate to $\beta$. Let’s call this variable $E$. The integral over $\beta,E$ is conditionally convergent if the $\beta$ integral is done first, imposing a constraint $\delta(E)$. [^14] The importance of doing the inverse Laplace transform integral over $\beta$ first was emphasized in [16]. The final answer is positive. To illustrate this in the simplest context, let’s arrange an action for $\beta$ and $E$ so that (4.12) arises from integrating out $E$:
@@ 5 Discussion
In this paper we followed a suggestion in [7] to use charged black holes as observers in de Sitter space, in order to understand the phase of the sphere partition function. Within a dilaton gravity reduction, we found the same phase $\mathrm{i}^{3}=-\mathrm{i}$ as in [7] for stable de Sitter black holes. For the unstable Nariai solution, we found trivial phase consistent with the results of [13, 12]. Our main technique to compute the phase was numerical, but we also studied the ultracold Nariai limit analytically. In this limit the black hole horizon and the de Sitter horizon become strongly fluctuating, governed by a $R^{2}$ two-dimensional gravity theory. We proposed a slightly different prescription from [7] for relating the Euclidean gravity path integral to a state-counting partition function. With this change, the density of states is real and positive, although this relies on the positivity of an integral with minus four zero modes: $\mathrm{i}^{4}=1$. It would be nice to have a more complete understanding of this factor.
@@
Clearly, many more checks must be performed to substantiate this. For example, one could study $G_{N}$ corrections to the existing theoretical evidence. It also seems important to understand the negative OTOC/signalling puzzle [48, 49, 50], the puzzle of the one-sided observer discussed in [8], and more generally, to analyze the potential instabilities of the sphere path integral in the presence of observers, aiming to develop a mathematically consistent prescription at the nonperturbative level. Ref. [9] already proposed a framework that, in principle, allows one to incorporate both perturbative and nonperturbative corrections. The proposal is that, when we focus on states defined in the presence of an observer, the Hartle–Hawking state becomes a natural state of maximal entropy in quantum gravity. Given the many lessons learned by studying near-extremal black holes in other settings, it might also be interesting to further study the ultracold limit, which is a natural state of minimal entropy.
@@ C.1 Method I: discrete mesh
$\mathcal{D}_{\text{mesh}}\Psi_{\text{mesh}}=\lambda\mathcal{M}_{\text{mesh}}\Psi_{\text{mesh}},\quad\quad\Psi_{\text{mesh}}\equiv\begin{pmatrix}\varphi_{1},&\cdots&\varphi_{n_{\text{mesh}}},&\omega_{1},&\cdots&\omega_{n_{\text{mesh}}}\end{pmatrix}^{\intercal}$, (C.8)
@@ C.2 Method II: fuzzy sphere regularization
With these ingredients, the fuzzy version of the differential operator $\mathcal{D}$ is now a $2(2S+1)^{2}$-by-$2(2S+1)^{2}$ dimensional matrix

### v2

@@ 1 Introduction
This puzzle was addressed by Maldacena in [41], motivated in part by the usefulness of including an observer in other studies of de Sitter space [10, 11, 17, 59]. [^2] It would be interesting to connect this proposal to the bulk vs. edge mode discussion in [8, 37]. The idea of [41] is that if an observer is included, all but two of the $D+1$ conformal Killing vectors become physical modes, because they move the observer. Or, in a different gauge, the observer adds D1D-1 new unstable modes that require Wick rotation. Either way, the phase now becomes $\mathrm{i}^{D+2-(D-1)}=\mathrm{i}^{3}=-\mathrm{i}$. [^3] Further work studied the phase for product manifolds [54, 32] and the Coleman-de Luccia instanton [33].
@@ 3.1 Setting up the computation
In the one-loop determinant, each negative mode gives a factor of $(-\mathrm{i})$, see appendix A. Our starting point in analyzing the phase is to define the measure so that it is real and positive if we Wick-rotate the entire conformal factor, or more generally one whole field [25]. In other words, we multiply the naive one-loop determinant by $(+\mathrm{i})^{n}$, where $n$ is the number of spacetime points. To evaluate the net phase, we would like to compare the number of negative eigenvalues of $\mathcal{D}$ to the number of spacetime points. In section 3.2, we use a pragmatic approach [^7] One could instead simply rotate one or the other of $\phi,\omega$ (or a linear combination) of them, but this leads to a complex fluctuation operator $\mathcal{D}$ that we found difficult to analyze. We also tried to define the regularized number of negative eigenvalues using a smooth cutoff in eigenvalue space, e.g. $\sum_{\lambda<0}e^{-\epsilon^{2}\lambda^{2}}$ and then dropping the $1/\epsilon$ term. However, because there are finite counterterms, this is not particularly unique, and in fact one finds that the answer depends on $Q$. that allows this comparison to be made. We discretize the manifold with a mesh, on which we study the spectrum of the quadratic fluctuations. For a finite mesh, the number of negative eigenvalues of $\mathcal{D}$ and the number of spacetime points are both finite and can be compared. We check this numerical answer with an analytical computation in the ultracold limit and an analytical understanding of how the phase changes at the special point $Q=1/4$.
@@ 3.3 A model for the lukewarm black holes
It is interesting to compare to the computation of the Euclidean Schwarzschild black hole in 4D flat space. This has a negative mode in the canonical ensemble [28, 6, 48] that can be explained by a similar integral, over the area-radius of the horizon: [^11] See appendix F.1 of [40] and more recently [38, 4] for a similar discussion of small black holes in AdS.
@@ 3.4 Ultracold Nariai
Here, $\omega_{\ell,m}$ are the spherical harmonic modes normalized using the physical metric $g$, not the metric $\widetilde{g}$. They are all unstable except the three $\ell=1$ zero modes, corresponding to the conformal Killing modes on the sphere (see (3.4)). [^13] The one-loop action in timelike Liouville is similar, but there the $\ell=0$ mode is positive [7]. The phase of the partition function comes from rotating all the unstable modes, each one gives a factor of $\mathrm{i}$. Due to the three missing zero modes, that leads to $\mathrm{i}^{3}$. The three zero modes also lead to a $q^{3}$ one-loop determinant. All together,
@@ 4 The phase of the state-counting partition function
In the previous sections, we studied the phase of the Euclidean gravity path integral. In this section, we compute the phase of the state-counting partition function. The difference is that the Euclidean gravity path integral is over real Euclidean geometries (up to contour rotations for convergence), whereas for the state-counting partition function we modify the integral in order to impose the Hamiltonian constraint $H=0$ [15, 43, 30, 19, 31, 41].
@@ 4.3 On the overall sign
We suggest a different prescription, based on the idea that the $\beta$ integral is supposed to impose the Hamiltonian constraint. To implement this, it is necessary to expose at least one further integral, over the energy variable conjugate to $\beta$. Let’s call this variable $E$. The integral over $\beta,E$ is conditionally convergent if the $\beta$ integral is done first, imposing a constraint $\delta(E)$. [^15] The importance of doing the inverse Laplace transform integral over $\beta$ first was emphasized in [43]. The final answer is positive. To illustrate this in the simplest context, let’s arrange an action for $\beta$ and $E$ so that (4.12) arises from integrating out $E$:
@@ 5 Discussion
In this paper we followed a suggestion in [41] to use charged black holes as observers in de Sitter space, in order to understand the phase of the sphere partition function. Within a dilaton gravity reduction, we found the same phase $\mathrm{i}^{3}=-\mathrm{i}$ as in [41] for stable de Sitter black holes. For the unstable Nariai solution, we found trivial phase consistent with the results of [32, 54]. Our main technique to compute the phase was numerical, but we also studied the ultracold Nariai limit analytically. In this limit the black hole horizon and the de Sitter horizon become strongly fluctuating, governed by a $R^{2}$ two-dimensional gravity theory. We proposed a slightly different prescription from [41] for relating the Euclidean gravity path integral to a state-counting partition function. With this change, the density of states is real and positive, although this relies on the positivity of an integral with minus four negative modes: $(-\mathrm{i})^{-4}=1$. It would be nice to have a more complete understanding of this factor.
@@
Clearly, many more checks must be performed to substantiate this. For example, one could study $G_{N}$ corrections to the existing theoretical evidence. It also seems important to understand the negative OTOC/signalling puzzle [23, 9, 2, 53], the puzzle of the one-sided observer discussed in [17], and more generally, to analyze the potential instabilities of the sphere path integral in the presence of observers, aiming to develop a mathematically consistent prescription at the nonperturbative level. Ref. [59] already proposed a framework that, in principle, allows one to incorporate both perturbative and nonperturbative corrections. The proposal is that, when we focus on states defined in the presence of an observer, the Hartle–Hawking state becomes a natural state of maximal entropy in quantum gravity. Given the many lessons learned by studying near-extremal black holes in other settings, it might also be interesting to further study the ultracold limit, which is a natural state of minimal entropy.
@@ C.1 Method I: discrete mesh
$\mathcal{D}_{\text{mesh}}\Psi_{\text{mesh}}=\lambda\mathcal{M}_{\text{mesh}}\Psi_{\text{mesh}},\quad\quad\Psi_{\text{mesh}}\equiv\begin{pmatrix}\varphi_{1},&\cdots&\varphi_{n_{\text{vertices}}},&\omega_{1},&\cdots&\omega_{n_{\text{vertices}}}\end{pmatrix}^{\intercal}$, (C.8)
@@ C.2 Method II: fuzzy sphere regularization
With these ingredients, the fuzzy version of the differential operator $\mathcal{D}$ is now an $2(2S+1)^{2}$-by-$2(2S+1)^{2}$ dimensional matrices